Andrew and Joseph took a trip to Vintage Vinyl Record Store at the store Andrew bought 6 records and 3 CDs for a total of $85.23. Joseph bought 4 records and 3 CDs for a total of $67.81


please help me write and solve a system of linear equations to find the cost of an individual record and an individual CD?
thank you!

Answer by Greenfinch(383)

I think you mean simultaneous equations.
6r + 3c = 8523
4r + 3c = 6781
so 2r = 1742
r = 871
substituting in (1) 5226 + 3c = 8523
3c = 3297
c = 1099
So a record costs $8.71 and a CD is $10.99